# Permissive Information-Flow Analysis for Large Language Models

### v2

@@ 5.1 Synthetic key-value dataset
Table 1: Results on the synthetic key-value dataset assuming a perfect recall retriever, which always retrieves the relevant items while also retrieving some irrelevant ones, with person ID as the key, and person SSN and DoB as values. $k$NN-LM prediction uses $\gamma=0.5$. We report mean and one standard deviation for macro-averaged precision and recall.
In this case, we assume a perfect recall retriever (all relevant documents are present), albeit lower precision (the rest of the documents out of the total limit of 14 are filled by adding irrelevant documents). For the introspection baseline, we use one-shot in-context learning [46] based on the first example in the dataset and use the remaining 63 questions for evaluation. Table 1 summarizes our results.
The prompt-based propagator achieves an exact match accuracy of over 85% and precision and recall of over 90%. That is, in more than 85% of the cases, the label search can identify the correct influential subcontexts out of a set of 16k possibilities without error. Furthermore, the prompt based label search significantly outperforms both the $k$NN-LM-based label search as well as the introspection baseline.
On the other hand, introspection performance improves significantly with the model scale. This result is consistent with the improved in-context learning ability observed in larger models [46]. Therefore, only the largest and most powerful models can be effectively used for introspection.
@@
Sensitivity of hyperparameters The perplexity tolerance $\lambda$ measures the acceptable loss in model utility when removing a subset of documents. The choice of $\lambda$ allows to trade-off model utility and the restrictiveness of the inferred label.
@@ 5.2 News article dataset
Moving towards a more realistic setup, we include an additional retriever component for the news article dataset that first retrieves relevant articles (either low or high integrity) from the dataset before response generation. In order to correctly understand the impact of the retriever, we compare a perfect retriever that is able to retrieve all relevant documents (similar to the synthetic key-value dataset) with a realistic retriever using cosine similarity in the embedding space computed by BGE-Large-EN [51].
Table 2 summarizes the results of the label propagator on the news article dataset. For this dataset, the label propagator achieves slightly worse performance on all metrics in comparison to the synthetic key-value dataset due to the lower quality of the dataset (automated generation of QA pairs by GPT-4 [32]). Furthermore, we see a reduction in the exact match accuracy of about $10\%$ when using a realistic retriever in contrast to a perfect retriever. This is due to the retriever sometimes failing to retrieve the relevant pieces of information, ultimately leading to a mismatch with the ground-truth label.
@@
Sensitivity of hyperparameters We highlight the sensitivity to changes in the perplexity threshold $\lambda$ on the news article dataset in Figure 8. Similar to the synthetic key-value dataset, we see similarities across the two model scales, and a higher tolerance of the larger model to lower values of $\lambda$. However, we see a significant rise in optimal perplexity tolerance in contrast to the synthetic dataset due to the higher difficulty of the responses.
@@ 5.3 LLM agent dataset
Model ROUGE-L$(y,y^{\star})$ ROUGE-L$(y^{\prime},y^{\star})$ ROUGE-L$(y^{\prime},y^{\star})-$ ROUGE-L$(y,y^{\star})$
@@
Table 4: Model output alignment with the dataset ground truth computed using ROUGE-L F-score on the LLM agent dataset. Utility computation is only applicable to the prompt-based label propagator. ROUGE-L$(y,y^{\star})$ refers to the alignment of the full context output (no regeneration required), while ROUGE-L$(y^{\prime},y^{\star})$ represents the alignment of the regenerated response after label improvement. Interestingly, we see that the regenerated response is more aligned with the ground truth response suggesting that in some cases label propagation can improve utility. ROUGE-L$(y^{\prime},y^{\star})-$ ROUGE-L$(y,y^{\star})$ compares the difference in alignment between the full context and the subcontext of the inferred label. When label improvement is possible, the difference in alignment is close to nil indicating that the regenerated response $y^{\prime}$ is highly aligned with the ground truth response $y^{\star}$. Otherwise, the difference in alignment is highly negative, indicating a significant drop in model alignment with $y^{\star}$.
@@
First, we focus on the 20 cases where label improvement is possible i.e., $L^{\star}\sqsubset\ell(C)$. We find that the label propagator improves the label in at least 17 of these cases.
In at most 2 of the cases, the LP is overly optimistic and returns a more permissive label $L^{\prime}$ than the ground truth label $L^{\star}$. Again, the safety property still holds in these cases because the system regenerates the output using the reduced context $C_{|L^{\prime}}$. This regeneration step introduces the possibility that the new output $y^{\prime}$ differs significantly compared to the output $y$ from the full context.
@@
In a robust retrieval augmentation setup, adding additional documents to the context should not negatively influence the model’s ability to answer a query since the model is free to ignore irrelevant inputs [39]. Thus, we expect that the ground truth completion $y^{\star}$ is always similar to the completion given the full context $y$. However, in almost all label propagation errors we observe, we find that the model is not able to answer the query with the full context. When controlling the model generation by using the ground truth target in the dataset instead of using the model-generated output, we see that our label propagator achieves perfect precision and recall.
@@ 6 Analysis and Discussion
Comparison to baselines The introspection-based label propagator achieves a noticeable improvement in performance when using a 70B parameter model compared to a 7B model. Furthermore, despite its simplicity and computational convenience, the introspection-based label propagator constitutes a strong baseline when considering simple lattices. We find a similar performance of the prompt-based label propagator for both model sizes suggesting that our approach could be implemented on smaller models, saving computational resources.
@@ 6.1 Use-cases
Our approach reveals its true benefits when the use of the generated content is not yet determined after the initial retrieval step. In such cases, having a too-restrictive label comes at the cost of limiting future uses of the generated output. For example, semantic caches [3] can be extended to store answers along with a sensitivity label, where more permissive labels facilitate broader reuse of the generated content. Likewise, Emails are often forwarded beyond their initial set of recipients, which is facilitated by using permissive labels.
@@ 6.2 Limitations
In the presence of an adversary, our label propagator assigns the lowest integrity label by design in order to avoid any harmful side effects such as cross-prompt injection [26]. However, this can potentially lead to a degradation of service attack as the adversary can add unreliable distracting information that makes it impossible for the system to improve the label, reducing the downstream utility of the model output.
@@ 6.3 Extensions
Efficiency improvements to Algorithm 1
Section 3.5 discusses the computational cost incurred to execute our label propagator in Algorithm 1. Despite an increase in the number of LLM calls for label propagation, the cost incurred per call can be drastically reduced utilizing common inference optimizations implemented in model serving backends [53, 21]. For instance, compute-bound prompt processing in Transformers can be amortized across calls sharing a common prompt prefix by reusing a KV cache. For a totally ordered lattice, appending retrieved documents at the end of the prompt sorted by their labels will result in no additional prompt processing costs in subsequent LLM calls because their prompts are obtained by peeling off documents at the end of the prompt.
We can also optimize the decoding phase of an LLM call by keeping a running calculation of the cumulative likelihood of the tokens decoded and stop decoding when the utility drops below a difference $\lambda$ of the utility of the full context.
Advanced influence estimation
The label propagation mechanism developed in Section 3 is based on estimating the influence of context elements on the model’s predictions. There are several alternatives techniques, such as Shapley value estimation, that can be used as drop-in replacements. See [4] for an overview of influence estimation techniques and [31, 8] for examples of their application to LLM contexts.
Another approach is to leverage the attention weights for influence estimation. Phenomena such as attention sinks [50] or lost-in-the-middle [25] makes identification based on simple attention scores difficult.
@@ 7 Related Work
Information-flow analysis and the use of flow labels has a long-standing history, see e.g., [9], with an early survey of language-based approaches [35]. The operations considered in many program analysis frameworks (e.g. [44, 35]) have well-defined semantics, and the rules for propagating the flow labels of each operator can be hard-coded. The situation is different when using ML models such as Transformers: due to their attention mechanism and autoregressive decoding, their output is tainted by all inputs. Moreover, their behavior is only implicitly defined as the solution of a loss-minimization problem. In our approach, we propagate labels through ML components under the constraint that the model’s loss does not increase.
@@
The literature distinguishes between static and dynamic approaches to information flow control [35, 5, 27]. The key difference is that dynamic approaches track the flow during program execution, whereas static approaches reason collectively about all possible executions without actually executing the program. Our approach relies on analyzing one specific context and is hence fundamentally dynamic.
@@
Secure multi-execution (SME) [10, 34] is an approach to dynamic information-flow analysis where programs are executed once per security level. The idea is that an execution that produces output for level $L$ receives data only from levels $L^{\prime}\sqsubseteq L$; and dummy values from levels $L^{\prime\prime}\sqsubseteq L$. Programs that are multi-executed are secure by design, and correct if the original program was secure. Our approach is related to SME in that we also call the LLM once per security level, and we propagate the most permissive label for which the original output is still "correct", in the sense that the drop in utility is bounded by $\lambda$.
Information-flow analysis in AI applications
@@
Several emerging approaches rely on information-flow tracking to obtain deterministic security guarantees for LLM-based applications, even if the underlying models are vulnerable to attacks: Wutschitz et al. [48] study how to leverage existing metadata in datastores used for retrieval augmentation to enforce confidentiality guarantees at inference time. Wu et al. [47] propose a system that uses information-flow tracking to prevent LLM-based applications from being compromised by untrusted information. [2] develop a monitor and policy language that can check execution traces of LLM applications for violations of different safety and information-flow properties. In contrast to these approaches, our work aims to dynamically identify labels of information that is propagated through LLM calls, which can be used to increase the permissiveness of system-level information-flow tracking.
@@
Related to the introspection baseline, Mireshghallah et al. [29] and Ghalebikesabi et al. [14] use Contextual Integrity theory to explore how well LLMs can be relied upon to discern which information is appropriate to use in a given context. Finally, Wallace et al. [45] train LLMs to distinguish between instructions in their context totally ordered in a hierarchy, teaching them to prioritize higher-privileged instructions over instructions that appear lower in the hierarchy.
@@ 8 Conclusion
9 Open Science
We will open-source our prototype implementation and the datasets used for its evaluation.
10 Ethics Considerations
We use a public dataset of fake and real news articles [33] as well as fully synthetic datasets for our evaluations. For instance, the synthetic key-value dataset contains entirely generated Social Security Numbers (SSNs) and dates of birth (DoB). Therefore, no private information from any individual is collected or used. Similarly, the LLM agent dataset is entirely synthetic and generated by us, ensuring that no real user interactions were recorded.
Additionally, we present a label propagation technique that can positively impact users of LLMs by providing hard safety guarantees, enhancing their reliability and security.
@@ Appendix A Improving performance on fine-grained lattices
In order to cover these cases when dealing with a complex lattice and a large number of relations, we propose a simple heuristic based on Shapley values that have been commonly used in the past [37, 31]. In particular, we compute the Shapley value for each of the different labels in the lattice and filter out any label combinations where labels below a particular threshold are present.

### v3

@@ 5.1 Synthetic key-value dataset
In this case, we assume a perfect recall retriever (all relevant documents are present), albeit lower precision (the rest of the documents out of the total limit of 14 are filled by adding irrelevant documents). For the introspection baseline, we use one-shot in-context learning (Wei et al., 2022) based on the first example in the dataset and use the remaining 63 questions for evaluation. Table 1 summarizes our results.
The prompt-based propagator achieves an exact match accuracy of over 85% and precision and recall of over 90%. That is, in more than 85% of the cases, the label search can identify the correct influential subcontexts out of a set of 16k possibilities without error. Furthermore, the prompt based label search significantly outperforms both the kkNN-LM-based label search as well as the introspection baseline.
On the other hand, introspection performance improves significantly with the model scale. This result is consistent with the improved in-context learning ability observed in larger models (Wei et al., 2022). Therefore, only the largest and most powerful models can be effectively used for introspection.
@@
Table 1: Results on the synthetic key-value dataset assuming a perfect recall retriever, which always retrieves the relevant items while also retrieving some irrelevant ones, with person ID as the key, and person SSN and DoB as values. kkNN-LM prediction uses $\gamma=0.5$. We report the mean and one standard deviation for macro-averaged precision and recall.
@@
Sensitivity of hyperparameters.
The perplexity tolerance $\lambda$ measures the acceptable loss in model utility when removing a subset of documents. The choice of $\lambda$ allows to trade-off model utility and the restrictiveness of the inferred label.
@@ 5.2 News article dataset
Moving towards a more realistic setup, we include an additional retriever component for the news article dataset that first retrieves relevant articles (either low or high integrity) from the dataset before response generation. In order to correctly understand the impact of the retriever, we compare a perfect retriever that is able to retrieve all relevant documents (similar to the synthetic key-value dataset) with a realistic retriever using cosine similarity in the embedding space computed by BGE-Large-EN (Xiao et al., 2023b).
Table 2 summarizes the results of the label propagator on the news article dataset. For this dataset, the label propagator achieves slightly worse performance on all metrics in comparison to the synthetic key-value dataset due to the lower quality of the dataset (automated generation of QA pairs by GPT-4 (OpenAI, 2023)). Furthermore, we see a reduction in the exact match accuracy of about $10\%$ when using a realistic retriever in contrast to a perfect retriever. This is due to the retriever sometimes failing to retrieve the relevant pieces of information, ultimately leading to a mismatch with the ground-truth label.
@@
Sensitivity of hyperparameters.
We highlight the sensitivity to changes in the perplexity threshold $\lambda$ on the news article dataset in Figure 8. Similar to the synthetic key-value dataset, we see similarities across the two model scales, and a higher tolerance of the larger model to lower values of $\lambda$. However, we see a significant rise in optimal perplexity tolerance in contrast to the synthetic dataset due to the higher difficulty of the responses.
@@ 5.3 LLM agent dataset
Model $(y,y^{\star}) ROUGE-L$$(y^{\prime},y^{\star})$ $(y^{\prime},y^{\star})- ROUGE-L$$(y,y^{\star})$
@@
Table 4: Model output alignment with the dataset ground truth computed using ROUGE-L F-score on the LLM agent dataset. Utility computation is only applicable to the prompt-based label propagator. ROUGE-L(y,y)($y^{\star}$) refers to the alignment of the full context output (no regeneration required), while ROUGE-L(y,y)($^{\prime},y^{\star}$) represents the alignment of the regenerated response after label improvement. Interestingly, we see that the regenerated response is more aligned with the ground truth response suggesting that in some cases label propagation can improve utility. $(y^{\prime},y^{\star})- ROUGE-L$$(y,y^{\star})$ compares the difference in alignment between the full context and the subcontext of the inferred label. When label improvement is possible, the difference in alignment is close to nil indicating that the regenerated response $y^{\prime}$ is highly aligned with the ground truth response $y^{\star}$. Otherwise, the difference in alignment is highly negative, indicating a significant drop in model alignment with $y^{\star}$.
@@
First, we focus on the 20 cases where label improvement is possible i.e., $L^{\star}\sqsubset\ell(C)$. We find that the label propagator improves the label in at least 17 of these cases. In at most 2 of the cases, the LP is overly optimistic and returns a more permissive label $L^{\prime}$ than the ground truth label $L^{\star}$. Again, the safety property still holds in these cases because the system regenerates the output using the reduced context $C_{|L^{\prime}}$. This regeneration step introduces the possibility that the new output $y^{\prime}$ differs significantly compared to the output $y$ from the full context.
@@
In a robust retrieval augmentation setup, adding additional documents to the context should not negatively influence the model’s ability to answer a query since the model is free to ignore irrelevant inputs (Shi et al., 2023). Thus, we expect that the ground truth completion $y^{\star}$ is always similar to the completion given the full context $y$. However, in almost all label propagation errors we observe, we find that the model is not able to answer the query with the full context. When controlling the model generation by using the ground truth target in the dataset instead of using the model-generated output, we see that our label propagator achieves perfect precision and recall.
@@ 6 Analysis and Discussion
Comparison to baselines.
The introspection-based label propagator achieves a noticeable improvement in performance when using a 70B parameter model compared to a 7B model. Furthermore, despite its simplicity and computational convenience, the introspection-based label propagator constitutes a strong baseline when considering simple lattices. We find a similar performance of the prompt-based label propagator for both model sizes suggesting that our approach could be implemented on smaller models, saving computational resources.
@@ 6.1 Use-cases
Our approach reveals its true benefits when the use of the generated content is not yet determined after the initial retrieval step. In such cases, having a too-restrictive label comes at the cost of limiting future uses of the generated output. For example, semantic caches (Bang, 2023) can be extended to store answers along with a sensitivity label, where more permissive labels facilitate broader reuse of the generated content. Likewise, Emails are often forwarded beyond their initial set of recipients, which is facilitated by using permissive labels.
@@ 6.2 Limitations
In the presence of an adversary, our label propagator assigns the lowest integrity label by design in order to avoid any harmful side effects such as cross-prompt injection (Liu et al., 2023). However, this can potentially lead to a degradation of service attack as the adversary can add unreliable distracting information that makes it impossible for the system to improve the label, reducing the downstream utility of the model output.
@@ 6.3 Extensions
Efficiency improvements to Algorithm 1.
Section 3.5 discusses the computational cost incurred to execute our label propagator in Algorithm 1. Despite an increase in the number of LLM calls for label propagation, the cost incurred per call can be drastically reduced utilizing common inference optimizations implemented in model serving backends (Zheng et al., 2023; Kwon et al., 2023). For instance, compute-bound prompt processing in Transformers can be amortized across calls sharing a common prompt prefix by reusing a KV cache. For a totally ordered lattice, appending retrieved documents at the end of the prompt sorted by their labels will result in no additional prompt processing costs in subsequent LLM calls because their prompts are obtained by peeling off documents at the end of the prompt.
We can also optimize the decoding phase of an LLM call by keeping a running calculation of the cumulative likelihood of the tokens decoded and stop decoding when the utility drops below a difference $\lambda$ of the utility of the full context. Liu et al. (2024a) also explored tricks to improve the efficiency of context attribution in LLMs.
Advanced influence estimation.
The label propagation mechanism developed in Section 3 is based on estimating the influence of context elements on the model’s predictions. There are several alternatives techniques, such as Shapley value estimation, that can be used as drop-in replacements. See Brophy et al. (2023) for an overview of influence estimation techniques and Nguyen & Wong (2023); Cohen-Wang et al. (2024) for examples of their application to LLM contexts.
Another approach is to leverage the attention weights for influence estimation. Phenomena such as attention sinks (Xiao et al., 2023a) or lost-in-the-middle (Liu et al., 2024b) makes identification based on simple attention scores difficult.
@@ 7 Related Work
Information-flow analysis and the use of flow labels has a long-standing history, see e.g., Denning (1976), with an early survey of language-based approaches (Sabelfeld & Myers, 2003). The operations considered in many program analysis frameworks (e.g. Volpano et al. (1996); Sabelfeld & Myers (2003)) have well-defined semantics, and the rules for propagating the flow labels of each operator can be hard-coded. The situation is different when using ML models such as Transformers: due to their attention mechanism and autoregressive decoding, their output is tainted by all inputs. Moreover, their behavior is only implicitly defined as the solution of a loss-minimization problem. In our approach, we propagate labels through ML components under the constraint that the model’s loss does not increase.
@@
The literature distinguishes between static and dynamic approaches to information flow control (Sabelfeld & Myers, 2003; Buiras et al., 2014; Mardziel et al., 2011). The key difference is that dynamic approaches track the flow during program execution, whereas static approaches reason collectively about all possible executions without actually executing the program. Our approach relies on analyzing one specific context and is hence fundamentally dynamic.
@@
Secure multi-execution (SME) (Devriese & Piessens, 2010; Rafnsson & Sabelfeld, 2016) is an approach to dynamic information-flow analysis where programs are executed once per security level. The idea is that an execution that produces output for level $L$ receives data only from levels $L^{\prime}\sqsubseteq L$; and dummy values from levels $L^{\prime\prime}\sqsubseteq L$. Programs that are multi-executed are secure by design, and correct if the original program was secure. Our approach is related to SME in that we also call the LLM once per security level, and we propagate the most permissive label for which the original output is still "correct", in the sense that the drop in utility is bounded by $\lambda$.
Information-flow analysis in AI applications.
@@
Several emerging approaches rely on information-flow tracking to obtain deterministic security guarantees for LLM-based applications, even if the underlying models are vulnerable to attacks: Wutschitz et al. (2023) study how to leverage existing metadata in datastores used for retrieval augmentation to enforce confidentiality guarantees at inference time. Wu et al. (2024) propose a system that uses information-flow tracking to prevent LLM-based applications from being compromised by untrusted information. Balunovic et al. (2024) develop a monitor and policy language that can check execution traces of LLM applications for violations of different safety and information-flow properties. An emergent line of work uses data and information-flow tracking to enforce deterministic security guarantees for the tools called by AI agents (Costa et al., 2025; Debenedetti et al., 2025; Zhong et al., 2025). In contrast to these approaches, our work aims to dynamically and soundly propagate labels through LLM calls, which can be used to increase the permissiveness of system-level information-flow tracking.
@@
Related to the introspection baseline, Mireshghallah et al. (2024) and Ghalebikesabi et al. (2024) use Contextual Integrity theory to explore how well LLMs can be relied upon to discern which information is appropriate to use in a given context. Finally, Wallace et al. (2024) train LLMs to distinguish between instructions in their context totally ordered in a hierarchy, teaching them to prioritize higher-privileged instructions over instructions that appear lower in the hierarchy.
@@ Appendix A Improving performance on fine-grained lattices
In order to cover these cases when dealing with a complex lattice and a large number of relations, we propose a simple heuristic based on Shapley values that have been commonly used in the past Shapley (1953); Nguyen & Wong (2023). In particular, we compute the Shapley value for each of the different labels in the lattice and filter out any label combinations where labels below a particular threshold are present.